OpenStudy (anonymous):

How can 1/3x − 2 =1/4x + 11 be set up as a system of equations?(

OpenStudy (anonymous):

A 3y − x = −6 4y − x = 44 B 3y + x = −6 4y + x = 44 C 3y + 3x = −6 4y + 4x = 44 D 3y − 3x = −6 4y − 4x = 44

OpenStudy (danjs):

yes, solve both equations for y, then set the x side equal

OpenStudy (danjs):

given y = 1/3 x - 2 y = 1/4 x + 11 there is a system in x and y, you can change it to standard form now

OpenStudy (anonymous):

Oh is it B?

OpenStudy (danjs):

how you get to that

OpenStudy (anonymous):

wonderful world of guessing?

OpenStudy (danjs):

They gfive you 1/3x − 2 =1/4x + 11 call both sides y, you have y=y, so if you break it into two equations , you have y = 1/3 x - 2 y = 1/4 x + 11 that is your sytem, rearrange both to the form of the answers ax+by = c

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

I got it its a :D